ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

16. Applications

1991 MS Classification: not given Kecs, Wilhelm, University of Petrosani, Romania The Solution of the Cauchy Problem for a Class of Operators which Describe the Vibrations of Viscoelastic Rods In the paper properties for the operators which describe the vibrations of viscoelastic rods are established. The solution of the Cauchy problem for this class of operators is given.